# Near-Discovery Observations of Interstellar Comet 3I/ATLAS with the NASA Infrared Telescope Facility

### v1

@@ Abstract
Interstellar Objects are comets and asteroids that formed around other stars but were ejected before they could accrete into exoplanets. They therefore represent a rare opportunity to compare the the building blocks of planets in the Solar System to those in other stellar systems. The third Interstellar Object, 3I/ATLAS, is the newest, brightest, potentially largest, and fastest member of this population. We report observations of 3I/ATLAS taken on 2025 July 3 and 4 with the NASA Infrared Telescope Facility just days after its discovery. In -band imaging with ’Opihi, we see no obvious lightcurve variability and derive a color of which is consistent in spectral slope to recent work by D. Z. Seligman et al. (2025) and C. Opitom et al. (2025). We obtained the first near-infrared (NIR) reflectance spectrum of 3I/ATLAS with SpeX. The visible color and NIR spectrum show a linear, red visible slope, a somewhat less red slope between 0.7 and 1.1 m, and a neutral or slightly blue slope at longer wavelengths. Challenges in modeling the reflectivity of 3I may indicate that this comet has a complex grain size distribution, grain compositions unlike Solar System comets, or both. Like 2I/Borisov, there are no obvious signatures of water ice in the coma of 3I/ATLAS. Observations closer to perihelion will help elucidate whether 3I has less water than anticipated or whether the Interstellar Objects might retain and release their ices somewhat differently than Solar System comets.
@@ 1 Introduction
The Interstellar Objects (ISOs) are small bodies analogous to our own comets and asteroids which formed around other stars, were subsequently ejected, and then were discovered passing through the Solar System on hyperbolic trajectories. The small bodies of the Solar System are its remnant planetsimals – the small portion of the planetesimals that did not get accreted to form the larger planets, get ejected, or fall into the Sun – and we expect that the same is true for the ISOs. Studying these objects individually and as a population therefore provides a means to probe the specifics of planet formation on a galactic scale. Two primary challenges encountered thus far have been the small number of objects discovered (before 2025, only two) and interpreting the the properties of both of those objects due to their difference from Solar System objects (see D. Jewitt & D. Z. Seligman 2023 for a recent review).
@@
The third ISO discovered, C/2025 N1 (ATLAS), now designated 3I/ATLAS, was found on 2025 July 1 by the ATLAS project’s (J. L. Tonry et al., 2018) telescope in Chile (see a review of its discovery circumstances in D. Z. Seligman et al. 2025). At discovery, the object was approximately in brightness and slightly diffuse, indicating that it was likely undergoing some kind of cometary mass loss. Given that 3I was discovered at opposition (), any potential tail would have been pointed away from the Earth and partially obscured by the foreground coma. 3I’s eccentricity () is significantly higher than the previous two ISOs and may be a good indication that it samples a different part of the galaxy (M. J. Hopkins et al., 2025) when taken with other aspects of its orbit. It is also intrinsically brighter than the previous two ISOs. Its bright absolute magnitude of cannot be directly converted into a diameter given the likely substantial coma contribution, but does constrain it to be at most km across (D. Z. Seligman et al., 2025).
@@
Spectral and photometric observations reported in D. Z. Seligman et al. (2025) show that 3I’s reflectance spectrum at visible wavelengths is approximately linear (without spectral curvature) and relatively red (), comparable to the surfaces of comets and D-type asteroids (see, e.g., F. E. DeMeo et al. 2009). Precovery observations suggest that the object may have been more neutral in color prior to its July 1 discovery (D. Z. Seligman et al., 2025). Observations from the Very Large Telescope (VLT) reported in C. Opitom et al. (2025) also show a red and linear visible slope (), while photometric observations from the Palomar 200” (B. T. Bolin et al., 2025) show a much more neutral visible reflectance slope ().
@@
Notably, despite higher spectral resolution and a larger aperture telescope, no gases were spotted in the vicinity of the comet in the VLT dataset (C. Opitom et al., 2025). This contrasts with the earliest observations of 2I/Borisov (see, e.g, A. Fitzsimmons et al. 2019; C. Opitom et al. 2019; T. Kareta et al. 2020) in which CN gas was detected rapidly with ground-based optical telescopes and then followed by other gases in quick succession (see, e.g., H. W. Lin et al. 2020; A. J. McKay et al. 2020). The absence of gas might be due to intrinsic differences between 3I and 2I, like the predicted water-dominated composition of 3I suggested by the modeling of M. J. Hopkins et al. (2025) connecting velocity and composition, but 3I/ATLAS was also discovered nearly au from the Sun, compared to approximately au for Borisov, and thus would be expected to be colder and less active. Although the larger heliocentric distance at which 3I was discovered might be an impediment to early detection of gas species, it might aid in the detection of water ice in its coma due to the longer time it takes ice to sublimate at those distances. Despite Borisov showing clear signs of water vapor production (see, e.g., Z. Xing et al. 2020), no clear signs of water ice were spotted in its coma (see, e.g., B. Yang et al. 2020; C.-H. Lee et al. 2020).
@@
In this paper, we present and analyze observations from SpeX, ‘Opihi, and MORIS on the NASA Infrared Telescope Facility taken within three days of 3I/ATLAS’s discovery to measure its reflectivity and search for signs of water ice. Comets, both interstellar and local, are time-variable on short and long timescales and thus the early nature of our observations should allow more direct comparison with other observations (e.g., those in the same activity state) in the discovery epoch than those taken later.
@@ 2 Observations
We observed 3I/ATLAS on UTC 2025 July 3 and 4 with SpeX (J. T. Rayner et al., 2003) on the NASA IRTF utilizing previously-scheduled (e.g., non-discretionary) time. SpeX was configured in the low-resolution ‘prism’ mode with a slit-width of ”, resulting in an effective resolution of approximately over the wavelength range. Given that the object was only expected to be ” in diameter, the 15”-long slit was used in a standard ABBA nodding pattern with the slit aligned along the parallactic angle (). Acquisition and guiding was accomplished with the visible-wavelength guider camera MORIS (A. A. S. Gulbis et al., 2010). We also utilized the wide-field acquisition camera ‘Opihi (E. Lee et al., 2022) with the Sloan filter on July 3 and alternating between the and filters on July 4 to obtain lightcurves and constrain 3I’s visible reflectivity. Given that time-sequence color photometry was being performed with ‘Opihi, filters were chosen for MORIS to try to maximize reliability of guiding as opposed to obtain visible colors or a lightcurve as is sometimes done in small body observations. Due to the object’s proximity to the galactic plane, acquisition of the object into the SpeX slit was challenging, and only frames in which we could be sure there was minimal contamination were reduced and analyzed further, corresponding to twenty frames on July 3 and fifty frames on July 4th. The on-chip integration time was 120 s per frame on both nights. Atmospheric conditions (i.e., seeing) were significantly better on July 4th, which, combined with the larger number of frames, led to better S/N for the spectrum that night. A local G3V standard star, HD164554, was observed at multiple airmasses both nights. The well-characterized solar analog Landoldt SA 107-684 was observed on July 3rd and BS4486 on July 4th.
@@
Simultaneous and imaging data were obtained using the IRTF’s ’Opihi facility camera (E. Lee et al., 2022). Attached to the IRTF’s 17-inch finder scope, the 20482048 CCD camera with 0.94 arcsecond pixels yields images with a 0.5 degree-on-a-side field of view. Sixty-second exposures were taken between dithers for the SpeX observations. Frames that were concurrent with telescope dithers were rejected, as were frames where the ISO was co-located on the sky with background objects of similar or greater brightness. On the night of July 3rd, the ’Opihi imaging yielded 129 band exposures spanning an interval of approximately 3 hours and 40 minutes. The following night of July 4th 72 band and 43 band images of sufficient quality were obtained over a timespan of 3 hours and 55 minutes.
@@
Reduction of the SpeX data followed the procedure described in J. P. Emery et al. (2011). Briefly, all frames were divided by a flat field (measured using an integrating sphere attached to Spex) and background frames were subtracted. Normally, one subtracts A-B pairs as a first-order sky emission removal, but the field was so crowded that many pairs contained stars in the background nod position. We therefore created median sky frames for each nod position each nigh using non-contaminated frames. The object position was located and summed within a 6-pixel-wide aperture. Standard stars were extracted similarly (though using normal A-B subtraction), and the object frames divided by star frames measured close in airmass. The resulting reflectance frames were normalized and averaged to compute the final spectrum each night. Wavelength calibration was applied using spectra of an Argon lamp and extracted using Spextool (M. C. Cushing et al., 2004). A portion of the July 4 dataset was reduced by a second co-author with the M. C. Cushing et al. (2004) package as a consistency check.
Reduction of the ’Opihi data was done using a combination of IRAF routines (D. Tody, 1986), and astropy python libraries (Astropy Collaboration et al., 2022). Aperture photometry was extracted using 4-arcsecond radius apertures on the object and a single Gaia DR3 (Gaia Collaboration et al., 2023) frame-reference star for each night. The extracted counts were subsequently corrected using full frame photometric zeropoint offsets obtained from fits to the band-specific Gaia catalog sources across each frame using astrometry.net software (D. Lang et al., 2010).
@@ 3.1 Visible Photometry
If we take the average apparent magnitude measured in Sloan over the three hours on July 3 and correct for geometric and heliocentric distances, we retrieve a phase-dependent absolute magnitude of . Although it is not clear what 3I’s actual phase curve looks like yet (and if its activity state is changing, it may be hard to determine for some time), typical cometary phase curves have linear coefficients of magnitudes per degree (R. Kokotanekova et al., 2017). Using this average phase coefficient for 3I, its true absolute magnitude would be . These results agree with measurements by other observers.
The following night and lightcurves were extraced from the 115 ’Opihi select frames taken on July 4 in which the target was very well separated from nearby stars. We derive an average color of from 10 images in each filter taken in a region of the sky that was relatively uncrowded by background stars of similar or greater brightness to 3I/ATLAS.
@@ 3.2 Near-Infrared Spectroscopy
We present the NIR reflectance spectrum of the 3I in Figure 2. The colors derived from ‘Opihi are also included, as well as the visible spectrum of 3I from D. Z. Seligman et al. (2025) for context. The NIR spectrum from July 3rd has significantly lower S/N, but agrees in spectral slope and shape, indicating no substantial change in the spectral properties of the coma over the two nights. As can be seen, the agreement between SpeX, ‘Opihi, and the visible spectrum from D. Z. Seligman et al. (2025) at the overlap wavelengths is quite good. The color derived from our best ‘Opihi frames, , is consistent with the visible slopes of D. Z. Seligman et al. (2025) and C. Opitom et al. (2025) but is significantly redder than the colors reported by B. T. Bolin et al. (2025). Although the short end of the NIR spectrum below is red and linear like most of the slopes reported for 3I at short wavelengths, the reflectance spectrum begins to decrease in slope beyond this point and approaches neutral reflectivity by . Beyond approximately , the spectrum is consistent with flat within errors. The amount of spectral curvature, especially considering the long-wavelength negative slopes, clearly differentiates the object from the D-type asteroids it appeared similar to with visible wavelength data only. While no asteroid type matches 3I’s spectrum well, the Double-Dip TNOs (N. Pinilla-Alonso et al., 2025) are quite similar in reflectivity. The D-type asteroids, the Double-Dip TNOs, and our comet model from Section 4 are also shown in Figure 2. However, at least some of the light we see is reflected from the object’s coma as opposed to from a solid surface, so the comparison between the asteroids, TNOs, or inactive other Solar System bodies may be of limited reliability.
@@ 4 Ice Abundance Modeling
In order to estimate the maximum allowable water ice fraction in the coma of 3I, we need to choose a model for the non-ice component (amorphous carbon is a typical choice for Solar System comets, whether the dust grains are porous), a physical state for the ice in the coma (physically separate or mixed into dust grains, crystalline vs. amorphous, etc.), and then a scattering mode for each kind of grains (e.g., using Mie theory or B. Hapke (2012) theory). We tested each of the combinations of models and optical constants used in T. Kareta et al. (2025), which were successfully validated on comets in stable activity and undergoing outbursts. Our approach was to model the NIR spectrum on its own, then take those models and see if they could reproduce the spectral behavior seen at visible wavelengths.

### v2

@@ Abstract
Interstellar Objects are comets and asteroids that formed around other stars but were ejected before they could accrete into exoplanets. They therefore represent a rare opportunity to compare the the building blocks of planets in the Solar System to those in other stellar systems. The third Interstellar Object, 3I/ATLAS, is the newest, brightest, potentially largest, and fastest member of this population. We report observations of 3I/ATLAS taken on 2025 July 3 and 4 with the NASA Infrared Telescope Facility just days after its discovery. In -band imaging with ’Opihi, we see no obvious lightcurve variability and derive a color of which is consistent in spectral slope to other near-discovery observations. We obtained the first near-infrared (NIR) reflectance spectrum of 3I/ATLAS with SpeX. The visible color and NIR spectrum show a linear, red visible slope, a somewhat less red slope between 0.7 and 1.1 m, and a neutral or slightly blue slope at longer wavelengths. Challenges in modeling the reflectivity of 3I may indicate that this comet has a complex grain size distribution, grain compositions unlike Solar System comets, or both. Like 2I/Borisov, there are no obvious signatures of water ice in the coma of 3I/ATLAS. Observations closer to perihelion will help elucidate whether 3I has less water than anticipated or whether the Interstellar Objects might retain and release their ices somewhat differently than Solar System comets.
@@ 1 Introduction
The Interstellar Objects (ISOs) are small bodies analogous to our own comets and asteroids which formed around other stars, were subsequently ejected, and then were discovered passing through the Solar System on hyperbolic trajectories. The small bodies of the Solar System are its remnant planetsimals – the small portion of the planetesimals that did not get accreted to form the larger planets, get ejected, or fall into the Sun – and we expect that the same is true for the ISOs. Studying these objects individually and as a population therefore provides a means to probe the specifics of planet formation on a galactic scale. Two primary challenges encountered thus far have been the small number of objects discovered (before 2025, only two) and interpreting the properties of both of those objects due to their difference from Solar System objects (see D. Jewitt & D. Z. Seligman 2023 for a recent review).
@@
The third ISO discovered, C/2025 N1 (ATLAS), now designated 3I/ATLAS, was found on 2025 July 1 by the ATLAS project’s (J. L. Tonry et al., 2018) telescope in Chile (see a review of its discovery circumstances in D. Z. Seligman et al. 2025). At discovery, the object was approximately in brightness and slightly diffuse, indicating that it was likely undergoing some kind of cometary mass loss. Given that 3I was discovered at opposition (phase angle ), any potential tail would have been pointed away from the Earth and partially obscured by the foreground coma. 3I’s eccentricity () is significantly higher than the previous two ISOs and may be a good indication that it samples a different part of the galaxy (M. J. Hopkins et al., 2025) when taken with other aspects of its orbit. It is also intrinsically brighter than the previous two ISOs. Its bright absolute magnitude of cannot be directly converted into a diameter given the likely substantial coma contribution, but does constrain it to be at most km across (D. Z. Seligman et al., 2025).
@@
Spectral and photometric observations reported in D. Z. Seligman et al. (2025) show that 3I’s reflectance spectrum at visible wavelengths is approximately linear (without spectral curvature) and relatively red (), comparable to the surfaces of comets and D-type asteroids (see, e.g., F. E. DeMeo et al. 2009). Precovery observations suggest that the object may have been more neutral in color prior to its July 1 discovery (D. Z. Seligman et al., 2025). Observations from the Very Large Telescope (VLT) reported in C. Opitom et al. (2025) also show a red and linear visible slope (), the Gran Telescopio Canarias (GTC) was used to measure (R. de la Fuente Marcos et al., 2025), and photometric observations from the Palomar 200” show similar slopes from multi-filter photometry (B. T. Bolin et al., 2025) and spectroscopy (M. Belyakov et al., 2025).
@@
Notably, despite higher spectral resolution and a larger aperture telescope, no gases were spotted in the vicinity of the comet in the VLT dataset (C. Opitom et al. 2025, see also A. Alvarez-Candal et al. 2025) or GTC dataset (R. de la Fuente Marcos et al., 2025). This contrasts with the earliest observations of 2I/Borisov (see, e.g, A. Fitzsimmons et al. 2019; C. Opitom et al. 2019; T. Kareta et al. 2020; J. de León et al. 2020) in which CN gas was detected rapidly with ground-based optical telescopes and then followed by other gases in quick succession (see, e.g., H. W. Lin et al. 2020; A. J. McKay et al. 2020). The absence of gas around the time of discovery might be due to intrinsic differences between 3I and 2I, such as the predicted water-dominated composition of 3I suggested by the modeling of M. J. Hopkins et al. (2025). 3I/ATLAS was, however, discovered nearly au from the Sun compared to approximately au for Borisov, and thus would be expected to be colder and less active early on. Although the larger heliocentric distance at which 3I was discovered might be an impediment to early detection of gas species, it might aid in the detection of water ice in its coma due to the longer time it takes ice to sublimate at those distances. Despite Borisov showing clear signs of water vapor production (see, e.g., Z. Xing et al. 2020), no clear signs of water ice were spotted in its coma (see, e.g., B. Yang et al. 2020; C.-H. Lee et al. 2020).
@@
In this paper, we present and analyze observations from SpeX and ‘Opihi on the NASA Infrared Telescope Facility (IRTF) taken within three days of 3I/ATLAS’s discovery to measure its reflectivity and search for signs of water ice. The early nature of our observations should allow more direct comparison with other observations (e.g., those in the same activity state) in the discovery epoch than those taken later.
@@ 2 Observations
We observed 3I/ATLAS on UTC 2025 July 3 and 4 with SpeX (J. T. Rayner et al. 2003, later upgraded in 2014) on the IRTF utilizing previously-scheduled (e.g., non-discretionary) time. On July 3 (July 4), 3I was () au away from the Sun, () au away from the Earth, and at a phase angle of (). SpeX was configured in the low-resolution ‘prism’ mode with a slit-width of ”, resulting in an effective resolution of approximately increasing to over the wavelength range. Given that the object was only expected to be ” in diameter, the 15”-long slit was used in a standard ABBA nodding pattern with the slit aligned along the parallactic angle (). Acquisition and guiding was accomplished with the visible-wavelength guider camera MORIS (A. A. S. Gulbis et al., 2010). We also utilized the wide-field acquisition camera ‘Opihi (E. Lee et al., 2022) with the Sloan filter on July 3 and alternating between the and filters on July 4 to obtain lightcurves and constrain 3I’s visible reflectivity. Given that time-sequence color photometry was being performed with ‘Opihi, filters were chosen for MORIS to try to maximize reliability of guiding as opposed to obtain visible colors or a lightcurve as is sometimes done in small body observations. Due to the object’s proximity to the galactic plane, acquisition of the object into the SpeX slit was challenging, and only frames in which we could be sure there was minimal contamination were reduced and analyzed further, corresponding to twenty frames on July 3 and fifty frames on July 4. The on-chip integration time was 120 s per frame on both nights. Atmospheric conditions (i.e., seeing) were significantly better on July 4, which, combined with the larger number of frames, led to better S/N for the spectrum that night. A local G3V standard star, HD164554, was observed at multiple airmasses both nights. The well-characterized solar analog Landoldt SA 107-684 was observed on July 3 and BS4486 on July 4.
@@
Simultaneous (July 3) and (July 4) imaging data were obtained using the IRTF’s ’Opihi facility camera (E. Lee et al., 2022). Attached to the IRTF’s 17-inch finder scope, the 20482048 CCD camera with 0.94 arcsecond pixels yields images with a square degrees field of view. (These large pixels made the extension of the target less obvious, but efforts to stack the ’Opihi and MORIS data will be pursued in more depth in a later paper.) Sixty-second exposures were taken between dithers for the SpeX observations. Frames that were concurrent with telescope dithers were rejected, as were frames where the ISO was co-located on the sky with background objects of similar or greater brightness. On the night of July 3, the ’Opihi imaging yielded 129 band exposures spanning an interval of approximately 3 hours and 40 minutes. The following night of July 4 72 band and 43 band images of sufficient quality were obtained over a timespan of 3 hours and 55 minutes.
@@
Reduction of the SpeX data followed the procedure described in J. P. Emery et al. (2011). Briefly, all frames were divided by a flat field (measured using an integrating sphere attached to Spex) and background frames were subtracted. Normally, one subtracts A-B pairs as a first-order sky emission removal, but the field was so crowded that many pairs contained stars in the background nod position. We therefore created median sky frames for each nod position each night using non-contaminated frames. The object position was located and summed within a 6-pixel-wide aperture. Standard stars were extracted similarly (though using normal A-B subtraction), and the object frames divided by star frames measured close in airmass. The resulting reflectance frames were normalized and averaged to compute the final spectrum each night. Wavelength calibration was applied using spectra of an Argon lamp and extracted using Spextool (M. C. Cushing et al., 2004). A portion of the July 4 dataset was reduced by a second co-author fully within the Spextool environment as a consistency check.
Reduction of the ’Opihi data was done using a combination of IRAF routines (D. Tody, 1986), and astropy python libraries (Astropy Collaboration et al., 2022). Per-frame astrometric solutions were computed using astrometry.net (D. Lang et al., 2010). Zero points for each frame were computed by matching tens of field stars to the PANSTARRS1 catalog (K. C. Chambers et al., 2016) using calviacat (M. Kelley & T. Lister, 2024), and thus reported magnitudes are within the PANSTARRS system. Aperture photometry was extracted using 4-pixel radius circular apertures on the object for all filters on both nights.
@@ 3.1 Visible Photometry
If we take the average apparent magnitude measured in Sloan over the three hours on July 3 and correct for geometric and heliocentric distances, we retrieve a phase-dependent absolute magnitude of . Although it is not clear what 3I’s actual phase curve looks like yet (and if its activity state is changing, it may be hard to determine for some time), typical cometary phase curves have linear coefficients of magnitudes per degree (R. Kokotanekova et al., 2017). Using this average phase coefficient for 3I, its true absolute magnitude would be . These results agree with measurements by other observers (see, e.g., a review of values in A. D. Feinstein et al. 2025).
The following night and lightcurves were extracted from the 115 ’Opihi select frames taken on July 4 in which the target was very well separated from nearby stars. We derive an average color of from the full observational set, including only images where 3I/ATLAS was not overlapping any identifiable background field star. As can be seen in Figure 2, this results in a visible slope quite similar to the D. Z. Seligman et al. (2025) spectrum, and thus the C. Opitom et al. (2025); B. T. Bolin et al. (2025); R. de la Fuente Marcos et al. (2025) spectra as well which report very similar slopes. B. T. Bolin et al. (2025) reports a color, which agrees with our measured color closer to less than one standard deviation. We do note that lightcurve effects and contamination from faint background sources may still be present, which could systematically alter the nightly-averaged color measurements we present here. Such effects are difficult to quantify given our data set, but we note that both the and data sets vary by magnitudes over the course of July 3 and July 4, respectively. We do not detect any clear trends in values, though this data was the most severely affected by background sources. Notably, R. de la Fuente Marcos et al. (2025) report a rotation period of hours for 3I/ATLAS based on near-discovery photometry (e.g., from near in time to ours). Our lightcurves strongly demonstrate limited ( magnitude) variability on short timescales, but this does not preclude somewhat larger variation like what they report ( magnitudes, R. de la Fuente Marcos et al. (2025)) so long as it is over this significantly longer timescale. While 3I/ATLAS has brightened somewhat since discovery making further attempts to measure its rotational period more challenging, it could be of interest to attempt to remeasure its nuclear lightcurve as it recedes from the Sun in 2026.
@@ 3.2 Near-Infrared Spectroscopy
We present the NIR reflectance spectrum of the 3I taken on July 4 in Figure 2. The color derived from ‘Opihi are also included, as well as the visible spectrum of 3I from D. Z. Seligman et al. (2025) for context. These and other datasets (described below) were all normalized at the wavelength of the Sloan (). The NIR spectrum from July 3rd has significantly lower S/N, but agrees in spectral slope and shape, indicating no substantial change in the spectral properties of the coma over the two nights. As can be seen, the agreement between SpeX, ‘Opihi, and the visible spectrum from D. Z. Seligman et al. (2025) at the overlap wavelengths is quite good. The color derived from our best ‘Opihi frames, , is consistent with the visible slopes of D. Z. Seligman et al. (2025) and C. Opitom et al. (2025) as well as the colors of B. T. Bolin et al. (2025) (). Although the short end of the NIR spectrum below is red and linear like most of the slopes reported for 3I at short wavelengths, the reflectance spectrum begins to decrease in slope beyond this point and approaches neutral reflectivity by . Beyond approximately , the spectrum is consistent with flat within errors. The amount of spectral curvature, especially considering the long-wavelength negative slopes, clearly differentiates the object from the D-type asteroids it appeared similar to with visible wavelength data only. While no asteroid type matches 3I’s spectrum well, the Double-Dip TNOs (N. Pinilla-Alonso et al., 2025) are quite similar in reflectivity. The D-type asteroids, the Double-Dip TNOs, and our comet model from Section 4 are also shown in Figure 2. However, at least some of the light we see is reflected from the object’s coma as opposed to from a solid surface, so the comparison between the asteroids, TNOs, or inactive other Solar System bodies may be of limited reliability.
@@ 4 Ice Abundance Modeling
In order to estimate the maximum allowable water ice fraction in the coma of 3I, we need to choose a model for the non-ice component (amorphous carbon is a typical choice for Solar System comets, whether the dust grains are porous), a physical state for the ice in the coma (physically separate or mixed into dust grains, crystalline vs. amorphous, etc.), and then a scattering mode for each kind of grains (e.g., using G. Mie (1908) theory or B. Hapke (2012) theory). We tested each of the combinations of models and optical constants used in T. Kareta et al. (2025), which were successfully validated on comets in stable activity and undergoing outbursts. Our approach was to model the NIR spectrum on its own, then take those models and see if they could reproduce the spectral behavior seen at visible wavelengths.